OpenStudy (anonymous):

Solve negative thirty-nine plus x is greater than or equal to negative forty-five for x.

OpenStudy (anonymous):

how do I do it

OpenStudy (anonymous):

help me please

OpenStudy (anonymous):

helpppppppppppppppppppppppppppppppppppppppppppppppppppppppppppppppppppppppppppppppppppppppppp

OpenStudy (anonymous):

iveeteee helllo

OpenStudy (ivettef365):

it is -39 + x >= -45

OpenStudy (anonymous):

yess

OpenStudy (ivettef365):

now is correct, the first thing you do is add 39 on both sides so you have -39 +x >= -45 +39 +39 --------------- x >= -6

OpenStudy (anonymous):

ok

OpenStudy (anonymous):